# Testing the Testers: Human-Driven Quality Assessment of Voice AI Testing Platforms

### v1

@@ 4.6.2 Test Scenarios
To select the final subset of scenarios and assign difficulties, we ran 3 simulations per provider and scenario using a calm and cooperative persona. This led to 270 simulations in total (3 providers with 3 simulations each on 30 candidate scenarios). The transcripts of these simulations were then evaluated by an LLM-as-a-judge prompt, which scored the Scenario Adherence metrics on a Likert scale [likert1932technique].
@@ 5.2.1 Metric Selection Criteria
  • CSAT (Customer Satisfaction Score) is widely used in industry [csat].
@@ 5.3.2 Human Participants
We recruited participants through Prolific [prolific2025] with strict eligibility criteria to ensure evaluator quality. Our sample consisted of $N_{e}=10$ participants per simulation. To prevent fatigue effects and ensure response diversity, we limited each participant to evaluating a maximum of three audio recordings.
@@ 5.4.1 Ground Truth Validation
Figure 4 presents the proportion of positive responses for each binary metric across all 600 human evaluations (60 simulations $\times$ 10 evaluators). The binary metrics demonstrated both strong agent performance and evaluator consensus, with positive rates ranging from 76.7% (Expected Outcome) to 95.0% (Avoid Repetition), highlighting the maturity of the chosen subject agent, provided by Sei  [sei] This variation indicates evaluators discriminated between different quality dimensions rather than uniformly rating all aspects positively.
@@ 5.4.2 Binary Metrics Performance
We employ four fundamental classification metrics [sokolova2009systematic]. Precision measures the proportion of positive predictions that were correct: $P=TP/(TP+FP)$, where $TP$ denotes true positives and $FP$ denotes false positives. Recall (sensitivity) measures the proportion of actual positive cases correctly identified: $R=TP/(TP+FN)$, where $FN$ denotes false negatives. The F1-score provides the harmonic mean of precision and recall: $F_{1}=2PR/(P+R)$, balancing both metrics to penalize extreme values [powers2011evaluation]. Accuracy measures overall correct predictions: $A=(TP+TN)/(TP+TN+FP+FN)$, where $TN$ denotes true negatives [fawcett2006introduction].
@@ 5.4.3 Continuous Metric Performance
We employ three complementary metrics to evaluate CSAT prediction performance [chai2014root]:
  • Mean Absolute Error (MAE) measures the average magnitude of prediction errors: $MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$, where $y_{i}$ represents the human rating and $\hat{y}_{i}$ the platform prediction. MAE provides an interpretable measure in the same units as the original scale [willmott2005advantages].
  • Root Mean Square Error (RMSE) penalizes larger errors more heavily: $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$, making it sensitive to outliers and particularly relevant for identifying platforms prone to severe mispredictions [hyndman2006another].
  • Pearson correlation coefficient measures the linear relationship between predicted and actual scores: $r=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y}_{i}-\bar{\hat{y}})}{\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\bar{\hat{y}})^{2}}}$, indicating whether platforms correctly capture relative satisfaction differences even if absolute calibration differs [benesty2009pearson].
These metrics collectively assess different aspects of prediction quality. MAE directly quantifies prediction accuracy in interpretable units (e.g., “predictions are off by 0.5 scale points on average”), crucial for operational decisions. RMSE identifies platforms that occasionally produce catastrophic errors—a critical consideration for customer service where severe mispredictions could miss escalation opportunities. Correlation reveals whether platforms understand the relative quality ordering of conversations, essential for comparative analytics and agent performance benchmarking [botchkarev2019performance].
@@ 5.4.4 Statistical Significance
Cochran’s Q test [cochran1950comparison] revealed highly significant differences among platforms ($Q(2)=62.85$, $p<0.001$), indicating the probability of observing such large differences by chance alone is less than $0.1\%$.
@@ 6.3 A final note on Platform Diversity and Market Evolution
Coval’s innovative application of autonomous vehicle testing methodologies to conversational AI, as noted by [techcrunch2025coval], represents a pioneering cross-domain approach that may yield future breakthroughs. The platform showed competitive performance in holistic assessment, achieving a 0.697 CSAT correlation, and demonstrated the highest precision scores across several metrics. Its simulation-first philosophy addresses a different aspect of the testing challenge, potentially offering advantages in scenario diversity and edge case discovery that our evaluation metrics may not fully capture.
@@ A.4 Code, transcripts and survey files
The code to reproduce the analysis, as well as the survey and response files, are all available in the accompanying repository https://github.com/evalionai/research. The conversational transcripts and audio files surveyed in this study are proprietary to Sei [sei] and we are not authorised to make them publicly available. For confidentiality reasons, only obfuscated versions of the transcripts are included in the repository. Access to the original transcripts and audio files may be granted for research purposes upon reasonable request, subject to client approval and confidentiality agreements. All methodological details are reported in the paper to ensure reproducibility. Researchers interested in data access may contact research@evalion.ai.

### v2

@@ 4.6.2 Test Scenarios
To select the final subset of scenarios and assign difficulties, we ran 3 simulations per provider and scenario using a calm and cooperative persona. This led to 270 simulations in total (3 providers with 3 simulations each on 30 candidate scenarios). The transcripts of these simulations were then evaluated by an LLM-as-a-judge prompt, which scored the Scenario Adherence metrics on a Likert scale [35].
@@ 5.2.1 Metric Selection Criteria
  • CSAT (Customer Satisfaction Score) is widely used in industry [29].
@@ 5.3.2 Human Participants
We recruited participants through Prolific [41] with strict eligibility criteria to ensure evaluator quality. Our sample consisted of $N_{e}=10$ participants per simulation. To prevent fatigue effects and ensure response diversity, we limited each participant to evaluating a maximum of three audio recordings.
@@ 5.4.1 Ground Truth Validation
Figure 4 presents the proportion of positive responses for each binary metric across all 600 human evaluations (60 simulations $\times$ 10 evaluators). The binary metrics demonstrated both strong agent performance and evaluator consensus, with positive rates ranging from 76.7% (Expected Outcome) to 95.0% (Avoid Repetition), highlighting the maturity of the chosen subject agent, provided by Sei  [44] This variation indicates evaluators discriminated between different quality dimensions rather than uniformly rating all aspects positively.
@@ 5.4.2 Binary Metrics Performance
We employ four fundamental classification metrics [48]. Precision measures the proportion of positive predictions that were correct: $P=TP/(TP+FP)$, where $TP$ denotes true positives and $FP$ denotes false positives. Recall (sensitivity) measures the proportion of actual positive cases correctly identified: $R=TP/(TP+FN)$, where $FN$ denotes false negatives. The F1-score provides the harmonic mean of precision and recall: $F_{1}=2PR/(P+R)$, balancing both metrics to penalize extreme values [40]. Accuracy measures overall correct predictions: $A=(TP+TN)/(TP+TN+FP+FN)$, where $TN$ denotes true negatives [19].
@@ 5.4.3 Continuous Metric Performance
We employ three complementary metrics to evaluate CSAT prediction performance [8]:
  • Mean Absolute Error (MAE) measures the average magnitude of prediction errors: $MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$, where $y_{i}$ represents the human rating and $\hat{y}_{i}$ the platform prediction. MAE provides an interpretable measure in the same units as the original scale [54].
  • Root Mean Square Error (RMSE) penalizes larger errors more heavily: $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$, making it sensitive to outliers and particularly relevant for identifying platforms prone to severe mispredictions [31].
  • Pearson correlation coefficient measures the linear relationship between predicted and actual scores: $r=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y}_{i}-\bar{\hat{y}})}{\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\bar{\hat{y}})^{2}}}$, indicating whether platforms correctly capture relative satisfaction differences even if absolute calibration differs [3].
These metrics collectively assess different aspects of prediction quality. MAE directly quantifies prediction accuracy in interpretable units (e.g., “predictions are off by 0.5 scale points on average”), crucial for operational decisions. RMSE identifies platforms that occasionally produce catastrophic errors—a critical consideration for customer service where severe mispredictions could miss escalation opportunities. Correlation reveals whether platforms understand the relative quality ordering of conversations, essential for comparative analytics and agent performance benchmarking [5].
@@ 5.4.4 Statistical Significance
Cochran’s Q test [11] revealed highly significant differences among platforms ($Q(2)=62.85$, $p<0.001$), indicating the probability of observing such large differences by chance alone is less than $0.1\%$.
@@ 6.3 A final note on Platform Diversity and Market Evolution
Coval’s innovative application of autonomous vehicle testing methodologies to conversational AI, as noted by [30], represents a pioneering cross-domain approach that may yield future breakthroughs. The platform showed competitive performance in holistic assessment, achieving a 0.697 CSAT correlation, and demonstrated the highest precision scores across several metrics. Its simulation-first philosophy addresses a different aspect of the testing challenge, potentially offering advantages in scenario diversity and edge case discovery that our evaluation metrics may not fully capture.
@@ A.4 Code, transcripts and survey files
The code to reproduce the analysis, as well as the survey and response files, are all available in the accompanying repository https://github.com/evalionai/research. The conversational transcripts and audio files surveyed in this study are proprietary to Sei [44] and we are not authorised to make them publicly available. For confidentiality reasons, only obfuscated versions of the transcripts are included in the repository. Access to the original transcripts and audio files may be granted for research purposes upon reasonable request, subject to client approval and confidentiality agreements. All methodological details are reported in the paper to ensure reproducibility. Researchers interested in data access may contact research@evalion.ai.